OpenStudy (anonymous):

what is the length of the focal chord? y²+6y-2x+13=0

OpenStudy (anonymous):

no idea for this one but i bet we could look it up

OpenStudy (anonymous):

i've searched everywhere for these and can't seem to find them on the graphing websites, do you know a good one?

OpenStudy (anonymous):

if you type in "parabola" it will give you the focus etc

OpenStudy (anonymous):

THANKS!!! so its 1 then? is the focal parameter the same as the length of the focal chord?

OpenStudy (anonymous):

no i think the length of the focal chord is the coefficient of the \(x\) term

OpenStudy (anonymous):

what would that be?

OpenStudy (anonymous):

\[y ^{2}+6y-2x+13=0\]\[y ^{2}+6y=2x-13\]complete the square \[y ^{2}+6y+9=2x-13+9\]\[(y+3)^{2}=2x-4\]standard form of a parabola\[(y+3)^{2}=2(x-2)\]vertex:(2,-3) 4p=2 4p is the length of the latus rectum... a focal chord parallel to the directrix.

OpenStudy (anonymous):

There are many focal chords though... I'm not sure which one your problem refers to... the most famous is the latus rectum... but a focal chord is a line segment that connects opposite sides of a parabola while passing through the focus.

OpenStudy (anonymous):

probably the most common one, what would that be?

OpenStudy (anonymous):

4p=2... the latus rectum is 2 units long for this parabola.

OpenStudy (anonymous):

It is the coefficient of the non-squared term in the equation above.

OpenStudy (anonymous):

so is 4p=2 the final answer?

OpenStudy (anonymous):

probably just 2

OpenStudy (anonymous):

thank you so much :) could one of you help me with another problem?

OpenStudy (anonymous):

post it new

OpenStudy (anonymous):

i will